# Renormalization group approach to graphene bilayers

### v1

@@ IV.4.1 Behaviour of the anomalous dimension $\eta_{k}^{B}$
Then, comparing the behaviour of $\eta_{k}^{B}$ and $\eta_{k}^{m}$ one also observes that the former stays longer in the harmonic regime ($\eta_{k}=0$) than the latter one. This relies on the very structure of $\eta_{B}^{k}$ which is such that, at the beginning of the flow, i.e. at large $\bar{c}_{k}$, one has
@@
IV.4.2 Behaviour of the Young modulus $\bar{Y}_{K}^{B}$
In Figs. 8, 9 and 10, we show the running Young modulus $\bar{Y}_{k}^{B}$ of a bilayer together with the quantity $2\,\bar{Y}_{k}^{m}$ corresponding to two uncoupled monolayers, for temperatures $T=10\,\text{K}$, $T=100\,\text{K}$ and $T=1000\,\text{K}$. Once again, we observe that at large RG-times the flow of the bilayer Young modulus $\bar{Y}_{k}^{B}$ converges towards $2\,\bar{Y}_{k}^{m}$, which itself approaches the value at the infrared fixed point of the flat phase, characterized by $\bar{\mu}_{*}^{m}\simeq 6.21$, $\bar{\lambda}_{*}^{m}\simeq-3.10$, and hence $2\,\bar{Y}_{*}^{m}\simeq 16.57$.
@@ IV.4.2 Behaviour of the Young modulus $\bar{Y}_{K}^{B}$
In this case, we see that $\bar{Y}_{k}^{B}$ exhibits a behaviour that differs drastically from that of a monolayer: it displays a very pronounced peak of order $650\%$ with respect to the monolayer value in the window between the harmonic-to-anharmonic crossover and the mechanical crossover. This phenomenon is again straightforward to analyze. To do so, let us consider the RG flow equation for $\bar{Y}_{k}^{B}$
@@
In Eq. (153) the dependence with respect to $\bar{c}_{k}$ appears both in $\eta_{k}$ and $H_{k}$. For large $\bar{c}_{k}$, $\eta_{k}$ is, as seen above, only slightly modified. However the behaviour of $H_{k}$ is given by
@@
and is thus of order unity, the nonlinear term in Eq. (153) becomes important as soon as $\bar{Y}_{K}$ becomes large. From these two behaviours one can thus expect a rapid change (a jump) of $\bar{Y}_{K}$ in a transient phase.
@@ V Conclusion
(i) As a first step, one can relax the assumption of a common field renormalization for both fields R and S and, more generally, of identical couplings for different kinds of monomials allowed by symmetry and entering in the action. In particular, it would make it possible to follow separately the renormalization of the symmetric and antisymmetric sectors and to test the robustness of the crossover scenario when the two modes acquire different anomalous dimensions. In the same spirit, the action can be enriched by terms involving higher powers of the fields R and S. Although simple power-counting and RG relevance arguments suggest that such contributions should remain subdominant for large-scale elasticity, including them explicitly would provide an a posteriori check of this expectation and a test of the stability of the critical exponents and crossover lengths.
@@
From a phenomenological perspective, it is natural to view the short-scale regime $\kappa_{\mathrm{eff}}\simeq\ell^{2}(\lambda+2\mu)/2$ and the long-scale regime $\kappa_{\mathrm{eff}}\simeq 2\kappa$ as the continuum counterparts of the distinct behaviours reported in recent experiments and simulations on few-layer graphene. The large apparent bending rigidity measured for bilayers in buckling and snap-through experiments [46] the pronounced softening and change of scaling under strong bending in trilayers and thicker stacks due to interlayer slip [50] and the observed size- and temperature-dependent effective stiffness in suspended membranes all point to a complex crossover structure [47, 48, 51]. In addition, atomistic Monte Carlo and molecular-dynamics simulations of bilayer and few-layer graphene also reveal a thickness-dependent flexural response and a crossover from correlated to uncorrelated layer fluctuations, further highlighting the role of interlayer coupling in the long-wavelength mechanical behaviour [28, 49]. Within our NPRG framework, this complexity is encoded in the flow of $\kappa_{\mathrm{eff}}(k)$ and the associated crossover scale $k_{c}$, which can be directly compared to the wave-vector windows effectively probed in these measurements and numerical studies.
@@ Appendix A Running anomalous dimension
where the threshold functions $\bar{L}_{abc}^{D+\alpha}$, $\bar{N}_{abc}^{D+\alpha}$ and $\bar{M}_{abc}^{D+\alpha}$ are given in the following section.
@@ B.1.1 Definitions
with $A_{D}=1/2^{D+1}\pi^{D/2}\Gamma(D/2)$, $P(\boldsymbol{q})=Z_{k}\,\boldsymbol{q}^{4}+R_{k}(\boldsymbol{q})$, $\widehat{\partial_{t}}=\partial_{t}R_{k}(\boldsymbol{q})\partial_{R_{k}}$.
Below we give the threshold functions $L_{a00}^{D+\alpha}$, $N_{a00}^{D+\alpha}$ and $M_{a00}^{D+\alpha}$ that are relevant to study the flat phase which is governed by massless flexurons modes. We thus put $m_{0}=0$. One has:
@@
We now set $x=\boldsymbol{q}^{2}$ so that $\text{d}^{D}q=2\,A_{D}\,x^{(D-2)/2}\text{d}x$ and apply $\widehat{\partial}_{t}=\partial_{t}R_{k}(\boldsymbol{q})\partial_{R_{k}}$. One gets:
$\left\{\begin{aligned} &L_{a00}^{D+\alpha}=\frac{1}{2}\int\text{d}x\,x^{{D+\alpha\over 2}-1}\frac{a\,\partial_{t}R_{k}(x)}{(Z_{k}\,x^{2}+R_{k}(x))^{a+1}}\\ \\ &\hskip-2.84544ptN_{a00}^{D+\alpha}=-\frac{1}{2}\int\text{d}x\,x^{{D+\alpha\over 2}-1}\left[\frac{\partial_{t}\partial_{x}P(x)}{(Z_{k}\,x^{2}+R_{k}(x))^{a}}\right.&\hskip 99.58464pt\left.-\,\frac{a\,\partial_{x}P(x)\,\partial_{t}R_{k}(x)}{(Z_{k}\,x^{2}+R_{k}(x))^{a+1}}\right]\\ \\ &M_{a00}^{D+\alpha}=-\frac{1}{2}\int\text{d}x\,x^{{D+\alpha\over 2}-1}\left[\frac{2\,\partial_{x}P(x)\,\partial_{t}\partial_{x}P(x)}{(Z_{k}\,x^{2}+R_{k}(x))^{a}}\right.\\ &\hskip 99.58464pt\left.-\frac{a\,(\partial_{x}P(x))^{2}\,\partial_{t}R_{k}(x)}{(Z_{k}\,x^{2}+R_{k}(x))^{a+1}}\right]\,.\end{aligned}\right$. (160)
@@ B.1.2 Dimensionless threshold functions
$R_{k}({\bf q})[y]=Z_{k}\,k^{4}\,y^{2}\,r_{k}(y)~~~~~\text{with}~~~~~y=\frac{x}{k^{2}}\,$, (161)
@@
$P(\boldsymbol{q})[y]=Z_{k}\,\boldsymbol{q}^{4}+R_{k}(\boldsymbol{q})=Z_{k}\,k^{4}\ y^{2}\big(1+r_{k}(y)\big)$ (162)
and compute all the quantities needed for the evaluations of the threshold functions Eq. (160). One has:
$\partial_{t}R_{k}(x)[y]=-Z_{k}\,k^{4}\big[\eta_{k}\,y^{2}r_{k}(y)+2y^{3}\partial_{y}r_{k}(y)\big]\\$ (163)
@@
$\widehat{\partial}_{t}\partial_{y}P(x)[y]=\partial_{y}\widehat{\partial}_{t}P(x)[y]=\partial_{y}\partial_{t}R_{k}(x)[y]$ (165)
@@ B.1.3 Litim (or $\Theta$) cut-off
We have chosen to use the Litim cut-off as cut-off function $R_{k}(\boldsymbol{q})$:
$R_{k}(\boldsymbol{q})[y]=Z_{k}(k^{4}-|\boldsymbol{q}|^{4})\Theta(k^{2}-|\boldsymbol{q}|^{2})=Z_{k}k^{4}r_{k}(y)$ (168)
@@ B.2.1 Definitions
with $P_{c_{k}}({\bf q})=\displaystyle\left(Z_{k}+\ell_{k}^{2}{{\lambda_{k}^{2}+2\mu_{k}^{2}}\over 4}\right){\bf q}^{4}+R_{k}({\bf q})$.
@@ B.2.2 Dimensionless threshold functions
$\hskip-22.76228pt\left\{\begin{aligned} &{}_{B}{}\bar{L}_{a00}^{D+\alpha}=-\frac{a}{2}\int\text{d}y\,y^{{{D+\alpha-2}\over 2}-2a}\left[\frac{\eta_{k}\,r_{k}(y)+2\,y\,\partial_{y}r_{k}(y)}{(1+\bar{c}_{k}+r_{k}(y))^{a+1}}\right]\\ &{}_{B}\bar{N}_{a00}^{D+\alpha}=\frac{1}{2}\int\text{d}y\,y^{{{D+\alpha}\over 2}-2a}\\ &\hskip 51.21504pt\left[\frac{2\,\eta_{k}\,r_{k}(y)+y(\eta_{k}+6)\partial_{y}r_{k}(y)+2\,y^{2}\partial^{2}_{y}r_{k}(y)}{(1+\bar{c}_{k}+r_{k}(y))^{a}}\right.\\ &\left.\hskip-5.69046pt-a\big(2(1+r_{k}(y)+\bar{c}_{k})\,+y\,\partial_{y}r_{k}(y)\big)\frac{\eta_{k}\,r_{k}(y)+2\,y\,\partial_{y}r_{k}(y)}{(1+\bar{c}_{k}+r_{k}(y))^{a+1}}\right]\end{aligned}\right$. (172)
@@ Appendix C Propagators of bilayer membranes
$\Delta$ $=4\,g_{2k}^{2}\,(\ell_{k}^{2}\boldsymbol{q}^{2}+\zeta_{k}^{2})^{2}+4\,g_{2k}\,\ell_{k}^{2}\boldsymbol{q}^{2}\,(\ell_{k}^{2}\boldsymbol{q}^{2}-\zeta_{k}^{2})\Big(3\,Z_{k}\,\boldsymbol{q}^{2}-\zeta_{k}^{2}\,(\lambda_{k}+2\mu_{k})\Big)+\ell_{k}^{4}\boldsymbol{q}^{4}\Big(3\,Z_{k}\,\boldsymbol{q}^{2}-\zeta_{k}^{2}\,(\lambda_{k}+2\mu_{k})\Big)^{2}$ (175)
$\ +3\,\ell_{k}^{2}\,R_{k}(\boldsymbol{q})\,\Big(4\,g_{2k}\,(\ell_{k}^{2}\boldsymbol{q}-\zeta_{k}^{2})+2\,\ell_{k}^{2}\boldsymbol{q}^{2}\,(3\,Z_{k}\,\boldsymbol{q}^{2}-\zeta_{k}^{2}\,(\lambda_{k}+2\mu_{k}))+3\,\ell_{k}^{2}\,R_{k}(\boldsymbol{q})\Big)\,.$
Appendix D Geometry of the bilayer: parallel surfaces

### v2

@@ IV.4.1 Behavior of the anomalous dimension $\eta_{k}^{B}$
Then, comparing the behavior of $\eta_{k}^{B}$ and $\eta_{k}^{m}$ one also observes that the former stays longer in the harmonic regime ($\eta_{k}=0$) than the latter one. This relies on the very structure of $\eta_{B}^{k}$ which is such that, at the beginning of the flow, i.e. at large $\bar{c}_{k}$, one has
@@
IV.4.2 Behavior of the Young modulus $\bar{Y}_{k}^{B}$
In Figs. 8-10, we show the running Young modulus $\bar{Y}_{k}^{B}$ of a bilayer together with the quantity $2\,\bar{Y}_{k}^{m}$ corresponding to two uncoupled monolayers, for temperatures $T=10\,\text{K}$, $T=100\,\text{K}$ and $T=1000\,\text{K}$. Once again, we observe that at large RG-times the flow of the bilayer Young modulus $\bar{Y}_{k}^{B}$ converges towards $2\,\bar{Y}_{k}^{m}$, which itself approaches the value at the infrared fixed point of the flat phase, characterized by $\bar{\mu}_{*}^{m}\simeq 6.21$, $\bar{\lambda}_{*}^{m}\simeq-3.10$, and hence $2\,\bar{Y}_{*}^{m}\simeq 16.57$.
@@ IV.4.2 Behavior of the Young modulus $\bar{Y}_{k}^{B}$
In this case, we see that $\bar{Y}_{k}^{B}$ exhibits a behavior that differs drastically from that of a monolayer: it displays a very pronounced peak of order $650\%$ with respect to the monolayer value in the window between the harmonic-to-anharmonic crossover and the mechanical crossover. This phenomenon is again straightforward to analyze. To do so, let us consider the RG flow equation for $\bar{Y}_{k}^{B}$
@@
In Eq. (142) the dependence with respect to $\bar{c}_{k}$ appears both in $\eta_{k}$ and $H_{k}$. For large $\bar{c}_{k}$, $\eta_{k}$ is, as seen above, only slightly modified. However the behavior of $H_{k}$ is given by
@@
and is thus of order unity, the nonlinear term in Eq. (142) becomes important as soon as $\bar{Y}_{k}$ becomes large. From these two behaviors one can thus expect a rapid change (a jump) of $\bar{Y}_{k}$ in a transient phase.
@@ V Conclusion
(i) As a first step, one can relax the assumption of a common field renormalization for both fields R and S and, more generally, of identical couplings for different kinds of monomials allowed by symmetry and entering in the running Gibbs free energy Eq.(70). In particular, it would make it possible to follow separately the renormalization of the symmetric and antisymmetric sectors and to test the robustness of the crossover scenario when the two modes acquire different anomalous dimensions. In the same spirit, the action can be enriched by terms involving higher powers of the fields R and S. Although simple power-counting and RG relevance arguments suggest that such contributions should remain subdominant for large-scale elasticity, including them explicitly would provide an a posteriori check of this expectation and a test of the stability of the critical exponents and crossover lengths.
@@
From a phenomenological perspective, it is natural to view the short-scale regime $\kappa_{\mathrm{eff}}\simeq\ell^{2}(\lambda+2\mu)/2$ and the long-scale regime $\kappa_{\mathrm{eff}}\simeq 2\kappa$ as the continuum counterparts of the distinct behaviors reported in recent experiments and simulations on few-layer graphene. The large apparent bending rigidity measured for bilayers in buckling and snap-through experiments [46] the pronounced softening and change of scaling under strong bending in trilayers and thicker stacks because of interlayer slip [50] and the observed size- and temperature-dependent effective stiffness in suspended membranes all point to a complex crossover structure [47, 48, 51]. In addition, atomistic Monte Carlo and molecular-dynamics simulations of bilayer and few-layer graphene also reveal a thickness-dependent flexural response and a crossover from correlated to uncorrelated layer fluctuations, further highlighting the role of interlayer coupling in the long-wavelength mechanical behavior [28, 49]. Within our NPRG framework, this complexity is encoded in the flow of $\kappa_{\mathrm{eff}}(k)$ and the associated crossover scale $k_{c}$, which can be directly compared to the wave-vector windows effectively probed in these measurements and numerical studies.
@@ Appendix A Running anomalous dimension
where the threshold functions $\bar{L}_{abc}^{D+\alpha}$, $\bar{N}_{abc}^{D+\alpha}$ and $\bar{M}_{abc}^{D+\alpha}$ are given in the following appendix.
@@ B.1.1 Definitions
with $A_{D}=1/2^{D+1}\pi^{D/2}\Gamma(D/2)$, $P(\boldsymbol{q})=Z_{k}\,\boldsymbol{q}^{4}+{\cal R}_{k}(\boldsymbol{q})$, $\widehat{\partial_{t}}=\partial_{t}{\cal R}_{k}(\boldsymbol{q})\partial_{{\cal R}_{k}}$.
Below we give the threshold functions $L_{a00}^{D+\alpha}$, $N_{a00}^{D+\alpha}$ and $M_{a00}^{D+\alpha}$ which are relevant to study the flat phase which is governed by massless flexurons modes. We thus put $m_{0}=0$. One has:
@@
We now set $x=\boldsymbol{q}^{2}$ so that $\text{d}^{D}q=2\,A_{D}\,x^{(D-2)/2}\text{d}x$ and apply $\widehat{\partial}_{t}=\partial_{t}{\cal R}_{k}(\boldsymbol{q})\partial_{{\cal R}_{k}}$. One gets:
$\left\{\begin{aligned} &L_{a00}^{D+\alpha}=\frac{1}{2}\int\text{d}x\,x^{{D+\alpha\over 2}-1}\frac{a\,\partial_{t}{\cal R}_{k}(x)}{(Z_{k}\,x^{2}+{\cal R}_{k}(x))^{a+1}}\\ &\hskip-2.84544ptN_{a00}^{D+\alpha}=-\frac{1}{2}\int\text{d}x\,x^{{D+\alpha\over 2}-1}\left[\frac{\partial_{t}\partial_{x}P(x)}{(Z_{k}\,x^{2}+{\cal R}_{k}(x))^{a}}\right.\\ &\hskip 99.58464pt\left.-\,\frac{a\,\partial_{x}P(x)\,\partial_{t}{\cal R}_{k}(x)}{(Z_{k}\,x^{2}+{\cal R}_{k}(x))^{a+1}}\right]\\ \\ &M_{a00}^{D+\alpha}=-\frac{1}{2}\int\text{d}x\,x^{{D+\alpha\over 2}-1}\left[\frac{2\,\partial_{x}P(x)\,\partial_{t}\partial_{x}P(x)}{(Z_{k}\,x^{2}+{\cal R}_{k}(x))^{a}}\right.\\ &\hskip 99.58464pt\left.-\frac{a\,(\partial_{x}P(x))^{2}\,\partial_{t}{\cal R}_{k}(x)}{(Z_{k}\,x^{2}+{\cal R}_{k}(x))^{a+1}}\right]\,.\end{aligned}\right$. (149)
@@ B.1.2 Dimensionless threshold functions
${\cal R}_{k}({\bf q})[y]=Z_{k}\,k^{4}\,y^{2}\,r_{k}(y)~~~~~\text{with}~~~~~y=\frac{x}{k^{2}}\,$, (150)
@@
$P(\boldsymbol{q})[y]=Z_{k}\,\boldsymbol{q}^{4}+{\cal R}_{k}(\boldsymbol{q})=Z_{k}\,k^{4}\ y^{2}\big(1+r_{k}(y)\big)$ (151)
and can compute all the quantities needed for the evaluations of the threshold functions Eq. (149). One has:
$\partial_{t}{\cal R}_{k}(x)[y]=-Z_{k}\,k^{4}\big[\eta_{k}\,y^{2}r_{k}(y)+2y^{3}\partial_{y}r_{k}(y)\big]\\$ (152)
@@
$\widehat{\partial}_{t}\partial_{y}P(x)[y]=\partial_{y}\widehat{\partial}_{t}P(x)[y]=\partial_{y}\partial_{t}{\cal R}_{k}(x)[y]$ (154)
@@ B.1.3 Litim (or $\Theta$) cut-off
We have chosen to use the Litim cut-off as cut-off function ${\cal R}_{k}(\boldsymbol{q})$:
${\cal R}_{k}(\boldsymbol{q})[y]=Z_{k}(k^{4}-|\boldsymbol{q}|^{4})\Theta(k^{2}-|\boldsymbol{q}|^{2})=Z_{k}k^{4}r_{k}(y)$ (157)
@@ B.2.1 Definitions
with $P_{c_{k}}({\bf q})=\displaystyle\left(Z_{k}+\ell_{k}^{2}{{\lambda_{k}^{2}+2\mu_{k}^{2}}\over 4}\right){\bf q}^{4}+{\cal R}_{k}({\bf q})$.
@@ B.2.2 Dimensionless threshold functions
$_{B}{}\bar{L}_{a00}^{D+\alpha}=-\frac{a}{2}\int\text{d}y\,y^{{{D+\alpha-2}\over 2}-2a}\left[\frac{\eta_{k}\,r_{k}(y)+2\,y\,\partial_{y}r_{k}(y)}{(1+\bar{c}_{k}+r_{k}(y))^{a+1}}\right]\\$ (161)
${}_{B}\bar{N}_{a00}^{D+\alpha}=\frac{1}{2}\int\text{d}y\,y^{{{D+\alpha}\over 2}-2a}$ (162)
$\left[\frac{2\,\eta_{k}\,r_{k}(y)+y(\eta_{k}+6)\partial_{y}r_{k}(y)+2\,y^{2}\partial^{2}_{y}r_{k}(y)}{(1+\bar{c}_{k}+r_{k}(y))^{a}}\right$.
$\left.\hskip-5.69046pt-a\big(2(1+r_{k}(y)+\bar{c}_{k})\,+y\,\partial_{y}r_{k}(y)\big)\frac{\eta_{k}\,r_{k}(y)+2\,y\,\partial_{y}r_{k}(y)}{(1+\bar{c}_{k}+r_{k}(y))^{a+1}}\right]$
@@ Appendix C Propagators of bilayer membranes
$\Delta(\boldsymbol{q})$ $=4\,g_{2k}^{2}\,(\ell_{k}^{2}\boldsymbol{q}^{2}+\zeta_{k}^{2})^{2}+4\,g_{2k}\,\ell_{k}^{2}\boldsymbol{q}^{2}\,(\ell_{k}^{2}\boldsymbol{q}^{2}-\zeta_{k}^{2})\Big(3\,Z_{k}\,\boldsymbol{q}^{2}-\zeta_{k}^{2}\,(\lambda_{k}+2\mu_{k})\Big)+\ell_{k}^{4}\boldsymbol{q}^{4}\Big(3\,Z_{k}\,\boldsymbol{q}^{2}-\zeta_{k}^{2}\,(\lambda_{k}+2\mu_{k})\Big)^{2}$ (165)
$\ +3\,\ell_{k}^{2}\,{\cal R}_{k}(\boldsymbol{q})\,\Big(4\,g_{2k}\,(\ell_{k}^{2}\boldsymbol{q}-\zeta_{k}^{2})+2\,\ell_{k}^{2}\boldsymbol{q}^{2}\,(3\,Z_{k}\,\boldsymbol{q}^{2}-\zeta_{k}^{2}\,(\lambda_{k}+2\mu_{k}))+3\,\ell_{k}^{2}\,{\cal R}_{k}(\boldsymbol{q})\Big)\,.$
Appendix D Low-wavevector expansion of the propagator $G_{k,3rs}({\bf q})$
One explains here the non-commutativity of the limits $\boldsymbol{q}^{2}\ell_{k}^{2}\ll\zeta_{k}^{2}$ and $g_{2k}\to 0$. One considers the propagator $G_{k,3rs}$ in the form
$G_{k,3rs}(\boldsymbol{q})\;=\;\frac{4\,\ell_{k}^{2}}{2\,g_{2k}\,(\ell_{k}^{2}\boldsymbol{q}^{2}+\zeta_{k}^{2})+\ell_{k}^{2}\!\left(5\,Z_{k}\,\boldsymbol{q}^{4}+\boldsymbol{q}^{2}\,\zeta_{k}^{2}\,(\lambda_{k}+2\mu_{k})+5\,{\cal R}_{k}(\boldsymbol{q})\right)-\sqrt{\Delta(\boldsymbol{q})}}\;,$ (166)
with $\Delta(\boldsymbol{q})$ the corresponding discriminant. The non-commutativity arises because the low-wavevector expansion, $\boldsymbol{q}^{2}\ell_{k}^{2}\ll\zeta_{k}^{2}$ is not uniform in $g_{2k}$, because of a cancellation between the explicit $\boldsymbol{q}^{2}\zeta_{k}^{2}(\lambda_{k}+2\mu_{k})$ term and the leading small-$\boldsymbol{q}$ contribution of $\sqrt{\Delta}$ when $g_{2k}\to 0$.
a. Limit $g_{2k}\to 0$ first.
At $g_{2k}=0$, the square root – taking ${\cal R}_{k}({\bf q})=0$ – admits the form:
$\sqrt{\Delta(\boldsymbol{q})}=\ell_{k}^{2}\,\boldsymbol{q}^{2}\,\zeta_{k}^{2}(\lambda_{k}+2\mu_{k})-3Z_{k}\,\ell_{k}^{2}\,\boldsymbol{q}^{4}\ .$ (167)
Inserting this expression into Eq.(166) the ${\cal O}(\boldsymbol{q}^{2})$ terms cancel in the denominator of $G_{k,3rs}(\boldsymbol{q})$ which is given by:
$\begin{array}[]{ll}\mathrm{Den}(\boldsymbol{q})&=8Z_{k}\,\ell_{k}^{2}\,\boldsymbol{q}^{4},\end{array}$ (168)
and therefore
$G_{k,3rs}(\boldsymbol{q})\;\xrightarrow[g_{2k}\to 0]{}\;\frac{4\ell_{k}^{2}}{8Z_{k}\,\ell_{k}^{2}\,\boldsymbol{q}^{4}}=\frac{1}{2Z_{k}\,\boldsymbol{q}^{4}}$ (169)
as expected.
b. Low-wavevector limit first, with fixed $g_{2k}>0$.
For fixed $g_{2k}>0$ and $\boldsymbol{q}^{2}\ell_{k}^{2}\ll\zeta_{k}^{2}$, the discriminant is dominated by its $g_{2k}^{2}$ contribution and
$\begin{array}[]{ll}\sqrt{\Delta(\boldsymbol{q})}&=2g_{2k}\,(\ell_{k}^{2}\boldsymbol{q}^{2}+\zeta_{k}^{2})+\ell_{k}^{2}\,\boldsymbol{q}^{2}\,\zeta_{k}^{2}(\lambda_{k}+2\mu_{k})\\ \\ &-\ell_{k}^{2}\,\boldsymbol{q}^{4}\bigl(3Z_{k}+2\ell_{k}^{2}(\lambda_{k}+2\mu_{k})\bigr)+{\cal O}(\boldsymbol{q}^{6}),\end{array}$ (170)
so that, in the denominator of (166), both the ${\cal O}(g_{2k})$ terms and the ${\cal O}(\boldsymbol{q}^{2})$ term cancel, leaving a leading ${\cal O}(\boldsymbol{q}^{4})$ contribution. One obtains
$\mathrm{Den}(\boldsymbol{q})\;\sim\;2\,\ell_{k}^{2}\,\boldsymbol{q}^{4}\bigl(4Z_{k}+\ell_{k}^{2}(\lambda_{k}+2\mu_{k})\bigr)$ (171)
hence
$\begin{array}[]{ll}G_{k,3rs}(\boldsymbol{q})&\;\sim\;\displaystyle\frac{4\ell_{k}^{2}}{2\,\ell_{k}^{2}\,\boldsymbol{q}^{4}\bigl(4Z_{k}+\ell_{k}^{2}(\lambda_{k}+2\mu_{k})\bigr)}\\ \\ &=\displaystyle\frac{1}{2\bigg(Z_{k}+{\ell_{k}^{2}(\lambda_{k}+2\mu_{k})\over 4}\bigg)\boldsymbol{q}^{4}}\end{array}$ (172)
that corresponds to Eq.(82).
Appendix E Geometry of the bilayer: parallel surfaces